OpenStudy (anonymous):

sum of $1000 was invested for 4 years, and the interest was compounded semiannually. If this sum amounted to $1389.08 in the given time, what was the interest rate? Please round your answer to two decimal places.

OpenStudy (anonymous):

Formula = P(1+R/N)^NT P = principle (how much you start out with) R = Interest Rate N = Number of times compounded a year T = Number of years 1000(1+R/2)^8 = 1389.08 divide by 1000 (1+R/2)^8 = 1.38908 take the eighth root of both sides 1+R/2 = 1.041936 subtract 1 then multiply by 2 R = .0839 or 8.39%